Form a partial differential equation by eliminating the function f from z = f(y/x)


1
Expert's answer
2022-03-02T05:52:40-0500

Given, "z = f\left(\dfrac{y}{x}\right)".


Differentiating partially with respect to "x", we get


"p=\dfrac{\partial z}{ \partial x}=f'\left(\dfrac{y}{x}\right)\cdot \dfrac{-y}{x^2}\\ \\

\therefore~f'\left(\dfrac{y}{x}\right)= -\dfrac{px^2}{y} \qquad\qquad~~~~~(1)"


Differentiating partially with respect to "y", we get

"q=\dfrac{\partial z}{ \partial y}=f'\left(\dfrac{y}{x}\right)\cdot \dfrac{1}{x}\\
f'\left(\dfrac{y}{x}\right) = qx \qquad\qquad\qquad~~~~~~~~(2)"


From (1) and (2), we get

"\begin{aligned}
-\dfrac{px^2}{y} &= qx\\
-px &= qy\\
xp + yq &= 0
\end{aligned}"

which is the required partial differential equation.